# A logic of co-valuations

### v1

@@ Abstract.
A co-valuation is, essentially, a minimal finite cover. We introduce a logic based on co-valuations, which play the role of valuations of free variables in classical first-order logic, and show that the fundamental tools of model theory – such as ultraproducts, compactness, and omitting types – can be developed in this framework. Using a recently discovered duality between certain countable posets and second-countable compact $T_{1}$ spaces, we prove that these spaces play the role analogous to countable universes in first-order logic. Thus, although no topology appears in the initial formulation, the logic of co-valuations turns out to be naturally suited for studying compact topological objects. Standard topological notions, such as connectedness and covering dimension, are easily expressible, and model-theoretic properties, including atomicity, can be effectively analyzed. The framework also interacts well with Fraïssé-type constructions.
@@ 1. Introduction
The first apporach is point-free topology, which ignores the points of the underlying space and focuses on algebraic structures derived from the topology, such as collections of open sets with operations like union, intersection, and closure. Early examples include Tarski’s work [13], which interprets intuitionistic logic via Heyting algebras of open sets; he also studied algebras of closed sets and, together with MacKinsey [11], closure algebras. Later, Henson et al. [10] investigated lattices of closed sets of topological spaces.
@@
Both strategies face inherent limitations. Algebraic structures derived from topologies are generally not axiomatizable, and two-sorted structures, where one sort represents the topology, are similarly resistant to axiomatization. As a result, the resulting logics often lack desirable properties: they may fail compactness, the downward Löwenheim-Skolem theorem does not hold, and so on.
One partial remedy is to restrict attention to subfamilies of topologies, such as topological bases. However, this is not a complete solution. A single space may admit multiple bases that are not elementarily equivalent. Expressibility is also limited: if one considers a sort for a basis, the downward Lǒwenheim-Skolem theorem holds, but only yields countable topological spaces. This severely restricts the ability to express fundamental topological properties such as connectedness or dimension – for instance, every countable metrizable space is homeomorphic to a subset of the rationals.
One can also focus on specific classes of spaces or alternative structures derived from a topological space. Bankston [1] considered regular compact spaces, which are equivalent, via Wallman duality, to disjunctive normal lattices. Nevertheless, the choice of a “right” basis remains problematic, and expressibility continues to depend heavily on the basis. On the other hand, Henson et al. [10] investigated also the ring $C(X)$ of real-valued functions on a topological space $X$. Similarly, Eagle et al. [5] associate a Hausdorff compact space $X$ with its $C^{*}$-algebra $C(X)$, and apply continuous logic. This is a promising research direction, however it seems restricted to spaces without any additional structure: what would be the $C^{*}$-algebra associated with a compact graph?
In this study, we take a different approach. The starting point is the notion of a co-valaution on a set $M$, which plays the role of a valuation in first-order logic – that is, an interpretation of free variables. Specifically, a co-valuation is a relation $v\subseteq X\times M$, where $X$ is a finite set (called a context) such that, for $x]_{v}=\{a\in M:(x,a)\in v\}$, the family $\{[x]_{v}:x\in X\}$ is a minimal cover of $M$. A fundamental relation between co-valuations is a consolidation: a covaluation $v\subseteq X\times Y$ is a consolidation of $w\subseteq Y\times M$ – and $w$ is a fragmentation of $v$ – if every set $x]_{v}$ is a union of elements of the form $y]_{w}$. The associated consolidation pattern is the relation $\geq\subseteq X\times Y$, defined by $x\geq y$ iff $x]_{v}\supseteq[y]_{w}$, which describes how the consolidation is carried out. A structure $(M,V)$ in the logic of co-valuations is an ordinary first-order structure $M$ together with a family $V$ of co-valuations on the universe of $M$ in contexts drawn from a fixed set of variables, satisfying certain additional requirements. Co-valuations in $V$ are called admissible.
@@
This logic behaves well with respect to ultraproducts: it satisfies Łoś’s theorem (Theorem 5.3), and therefore enjoys compactness (Corollary 5.2). Moreover, with an aid of a duality recently discovered by Bartoš et al. [2] between certain countable posets, called $\omega$-posets, and second-countable compact $T_{1}$ spaces, for every structure, one can construct an elementarily equivalent one whose admissible co-valuations generate a second-countable compact $T_{1}$ topology. We call such structures compact $\omega$-structures. Thus, a topologically well-behaved downward Löwneheim-Skolem theorem holds (Theorem 5.3 ), even though no topology is present in the original formulation. Furthemore, the omitting types theorem can be proved (Theorem 5.6), using Baire-category methods in the Polish space $\rm{Mod}(L)$ of all compact $\omega$-structures in signature $L$. This leads to a natural notion of atomicity for compact $\omega$-structures (Theorem 5.12).
@@
One can show that every compact $T_{1}$ space equipped with closed relations can be naturally represented as a structure in our logic of co-valuations (Corollary 4.13). Since admissible co-valuations form only a basis for the underlying topology, not all topological properties are expressible in the logic. Nevertheless, notions such as connectedeness, compactness, perfectness, or covering dimension are expressible; some others, such as metrizability or chainability follow from omitting a suitable family of types (Section 4.4).
@@
The fact that admissible co-valuations give only a basis is a limitation. However, for many classical topological objects there is a canonical choice of admissible co-valuations. Specifically, this occurs for second-countable compact $T_{1}$ spaces (possibly with additional relations) that arise as Fraïssé limits in a variant of Fraïssé theory suitable for this setting (Section 6). When admissible co-valuations are obtained via a sufficiently nice ($L$-faithful) Fraïssé sequence in an appropriate category, the resulting structure is an atomic model of its theory (Theorem 6.6), and any two such structures are isomorphic (Theorem 6.5). This applies, for example, to the arc, the ordered arc, and the pseudo-arc. Thus, for compact objects arising as Fraïssé limits, there is a canonical way to select the “correct” corresponding structure in the logic of co-valuations.
Finally, we remark that the present study may also be viewed as a contribution to a more abstract theory of valuations in logic. What, in fact, are the common structural features of the ordinary valuation and of the co-valuation introduced in this paper? The interpretation of free variables – arguably one of the most basic semantic operations in logic – has largely been treated as conceptually straightforward and, to the best of our knowledge, has never been examined in its own right. Our results suggest that this operation has a richer mathematical character than is usually assumed. A systematic investigation of the mathematical nature of valuation, and of possible generalizations such as co-valuation, therefore appears to be a promising direction for future research.
@@
The paper is organized as follows. Section 3 provides a brief overview of the Bartoš-Bice-Vignati duality between $\omega$-posets and second- countable compact $T_{1}$ spaces. In Section 4, we introduce the logic of co-valuations, define compact structures, and examine several basic topological properties that are expressible in this logic. Section 5 develops the model theory: ultraproducts, Łoś’s theorem, the downward Löwenheim-Skolem theorem, omitting types, and atomic models. Section 6 presents a variant of Fraïssé theory adapted to the logic of co-valuations, and discusses atomicity of Fraïssé limits along with several examples.
@@ 2. Relations
$xR^{\prime}y,\,x^{\prime}Ry\rightarrow x=x^{\prime}.$
@@
We say that $A$ is $R$-refined by $B\subseteq Y$ (or $B$ is refined by $A$, when $R$ is clear from the context) if $A^{R}\supseteq B$. In the case that the ordering relation of a poset $\mathbb{P}$ is not explicitely specified, we will refer to it by $\leq$ (or by $\geq$). For $A,B\subseteq\mathbb{P}$, we say that $A$ is an $\leq$-up-set (or up-set) if $A^{\leq}=A$,
@@ 3. Constructing compact spaces from posets
In this section, we present a brief overview of the Bartoš–Bice–Vignati duality connecting certain countable posets and second-countable compact $T_{1}$ spaces in the spirit of Stone duality. It has been developed in [2], to which the reader is referred for a detailed exposition. Here we present several proofs to illustrate the roles played by the main concepts of this theory.
We begin with some motivation. Let $X$ be a second-countable compact $T_{1}$ space, and let $\mathbb{P}_{n}$, $n\in\mathbb{N}$, be a sequence of finite open covers of $X$ such that every $\mathbb{P}_{n}$ is refined by $\mathbb{P}_{n+1}$, and every open cover of $X$ is refined by some $\mathbb{P}_{n}$. In particular, $\bigcup\mathbb{P}_{n}$ is a basis of $X$ (see Proposition 3.6). For $T_{2}$, i.e., metrizable, compact spaces, the construction of such a sequence, using the Lebesgue number of a finite cover, is straightfoward; the general case is also elementary (see [2, Proposition 1.11]). Let the disjoint union $\mathbb{P}=\bigsqcup\mathbb{P}_{n}$ be ordered by inclusion.
@@
As a matter of fact, we will work with $\omega$-posets of a particularly simple form, where the ordering relation is determined by relations between consequtive levels. Let $\mathbb{P}_{n}$, $n\in\mathbb{N}$, be finite sets, and let $\geq_{n}\subseteq\mathbb{P}_{n}\times\mathbb{P}_{n+1}$, $n\in\mathbb{N}$, be co-surjective relations. We denote by $\geq^{m}_{n}$, where $m<n$, appropriate compositions of $\geq_{n}$. Let $\mathbb{P}=\bigsqcup\mathbb{P}_{n}$, and let $\geq$ be the reflexive and transitive closure of $\bigsqcup\geq_{n}$. It is straightforward to verify that the resulting poset $(\mathbb{P},\geq)$ is an $\omega$-poset, referred to as an $\omega$-chain $(\mathbb{P}_{n},\geq_{n})$. For an $\omega$-chain $(\mathbb{P},\geq)$, $\mathbb{P}_{n}$ always denote the levels of $\mathbb{P}$, and $\geq_{n}=\geq\upharpoonright\mathbb{P}_{n}\times\mathbb{P}_{n+1}$, $n\in\mathbb{N}$. We call $(\mathbb{P}_{n},\geq_{n})$ level-injective if all $\geq_{n}$ are co-injective.
@@
Let $\mathbb{P}_{n}$, $n\in\mathbb{N}$, be minimal covers of a set $X$ such that each $\mathbb{P}_{n}$ consolidates $\mathbb{P}_{n+1}$. Then $(\bigsqcup\mathbb{P}_{n},\supseteq)$ is the level-injective $\omega$-chain $(\mathbb{P}_{n},\supseteq\upharpoonright\mathbb{P}_{n}\times\mathbb{P}_{n+1})$.
@@
Let $\supseteq_{n}=\supseteq\upharpoonright\mathbb{P}_{n}\times\mathbb{P}_{n+1}$, and, for $m<n$, denote by $\supseteq^{m}_{n}$ appropriate compositions of $\supseteq_{n}$. We claim that if $\supseteq_{n}^{0}$ is the transitive closure of $\bigcup_{m\leq n}\supseteq_{m}$, then $\supseteq_{n+1}^{0}$ is the transitive closure of $\bigcup_{m\leq n+1}\supseteq_{m}$. Fix $m\in\mathbb{N}$ with $m<n$, and $p\in\mathbb{P}_{m}$, $q\in\mathbb{P}_{n+1}$ with $p\supseteq q$. By minimality of the cover $\mathbb{P}_{n+1}$, there is $x\in q\setminus\bigcup\{q^{\prime}\in\mathbb{P}_{n+1}:q^{\prime}\neq q\}$. Because $\mathbb{P}_{m}$ consolidates $\mathbb{P}_{n}$, there is $q^{\prime}\in\mathbb{P}_{n}$ such that $x\in q^{\prime}\subseteq p$. But then we must have $q\subseteq q^{\prime}$, which immediately implies the claim. ∎
@@
For each $x\in X$, let $d_{x}=\{a\in A\cup B:x\in a\}$. As $A$ and $B$ are finite, so is $D=\{d_{x}:x\in X\}$. For each $d\in D$, choose some $x_{d}$ such that $d=d_{x_{d}}$ and denote the set of all the other chosen points by
@@
Also note that if $y\in a\in A$ then $y\neq x_{d}$ for any $d\neq d_{y}$ (because $y=x_{d}$ implies $d_{y}=d_{x_{d}}=d$) so $y\in d_{y}\setminus f_{d_{y}}\subseteq a$. This shows that $a=\bigcup\{c\in C:c\subseteq a\}$, for all $a\in A$, i.e., $C$ fragments $A$. For the same reason, $C$ fragments $B$. ∎
@@
Fix an open $U$, and $x\in U$. Let $C$ be an open cover of $X$ extending $\{U\}$, and such that $x\not\in\bigcup C\setminus\{U\}$. Fix $C^{\prime}\in\mathcal{C}$ that refines $C$, and let $V\in C^{\prime}$ be such that $x\in V$. Then we must have $V\subseteq U$. ∎
@@
For $\omega$-chains without minimal elements, e.g., level-injective $\omega$-chains, caps can be equivalently defined as sets refined by a level.
@@
To prove (1), we use Zorn’s lemma. Let $S_{i}$, $i\in I$, be an infinite chain (under inclusion) of selectors. We show that $S=\bigcap_{i}S_{i}$ is a selector. Let $C$ be a cap. We can assume that $C$ is finite, so there is $c\in C$ such that for $S_{i}\cap C=\{c\}$ for almost all $i\in I$. But $S_{i}$ form a chain, so, actually, $S_{i}\cap C=\{c\}$ for all $i\in I$.
@@
A basis $\mathbb{P}$ of a topological space $X$ is called a cap-basis if every cover of $X$ consisting of sets from $\mathbb{P}$ is a cap in $\mathbb{P}$ ordered by inclusion.
@@
We show that $\tau_{\mathbb{P}}$ is a compact topology. Without loss of generality, we can assume that $\mathsf{S}$$\mathbb{P}$, and so $\mathbb{P}$, is infinite. Fix a cap $C$, and a band $B$ that refines $C$. Clearly, $B^{\in}$ is a band in $\mathbb{P}^{\in}$, i.e., $C^{\in}$ is a cap. Fix $S\in{\mathsf{S}}\mathbb{P}$. As $S$ is infinite, $B$ is finite, and $\mathbb{P}$ is an $\omega$-poset, there must exist $s\in S$, and $b\in B$ such that $s\leq b$. Because $B$ refines $C$, there is $c\in C$ such that $b\leq c$. But $S$ is an up-set, so $c\in S$, i.e., $S\in c^{\in}$. As $S$ was arbitrary, we get that $C^{\in}$ covers $\mathsf{S}$$\mathbb{P}$.
Suppose now that $C\subseteq\mathbb{P}$ is not a cap. In other words, $\mathbb{P}\setminus C$ is a selector, so it contains a minimal selector $S$. But then $c\not\in S$, i.e., $S\not\in c^{\in}$ for every $c\in C$, and $C_{S}$ does not cover $\mathsf{S}$$\mathbb{P}$. Therefore covers of $X$ by sets from $\mathbb{P}^{\in}$ are given by caps in $\mathbb{P}$. In particular, $\mathbb{P}^{\in}$ is a cap-basis, and $\tau_{\mathbb{P}}$ is compact. ∎
@@
If $\mathbb{P}$ is a level-injective $\omega$-chain, then every $\mathbb{P}_{n}^{\in}$ is a minimal cover of $\mathsf{S}$$\mathbb{P}$. In particular, $p^{\in}\neq\emptyset$, for $p\in\mathbb{P}$.
@@
Note that $x\in p$ iff $p\in x^{\in}$, so the homeomorphism $x\mapsto x^{\in}$ maps $p$ onto $p^{\in}$.
@@
Let $\mathbb{P}$ be a level-injective $\omega$-chain. Then $x\geq y$ iff $x^{\in}\supseteq y^{\in}$, for $x\in\mathbb{P}_{m}$, $y\in\mathbb{P}_{n}$, $m\leq n$.

### v2

@@ Abstract.
A co-valuation is, essentially, a minimal finite cover. We introduce a logic based on co-valuations, which play the role of valuations of free variables in classical first-order logic, and show that the fundamental tools of model theory – such as ultraproducts, compactness, and omitting types – can be developed in this setup. Using a recently discovered duality between certain countable posets and second-countable compact $T_{1}$ spaces, we show that these spaces are counterparts of countable universes in first-order logic. Thus, although no topology appears in the initial formulation, the logic of co-valuations turns out to be naturally suited for studying compact topological objects. Standard topological notions, such as connectedness and covering dimension, are easily expressible, and model-theoretic properties, such as atomicity, can be effectively analyzed. The framework also interacts well with Fraïssé-type constructions.
@@ 1. Introduction
The first apporach is point-free topology, which ignores the points of the underlying space and focuses on algebraic structures derived from the topology, such as collections of open sets with operations like union, intersection, and closure. Early examples include Tarski’s work [13], which interprets intuitionistic logic via Heyting algebras of open sets; Tarski also studied algebras of closed sets and, together with MacKinsey [11], closure algebras. Later, Henson et al. [10] investigated lattices of closed sets of topological spaces.
@@
Both strategies face inherent limitations. Algebraic structures derived from topologies are generally not axiomatizable, and two-sorted structures, where one sort represents the topology, are similarly resistant to axiomatization. As a consequence, the resulting logics often lack desirable properties: they may fail compactness, the downward Löwenheim-Skolem theorem does not hold, etc.
One partial remedy is to restrict attention to subfamilies of topologies, such as topological bases. However, this is not a complete solution. A single space may admit multiple bases that are not elementarily equivalent. Expressibility is also limited: if one considers a sort for a basis, the downward Löwenheim-Skolem theorem holds, but only yields countable topological spaces. This severely restricts the ability to express fundamental topological properties such as connectedness or dimension – for instance, every countable metrizable space is homeomorphic to a subset of the rational numbers.
One can also focus on specific classes of spaces or alternative structures derived from a topological space. Bankston [1] considered regular compact spaces, which are equivalent, via Wallman duality [14], to disjunctive normal lattices. Nevertheless, the choice of a “right” basis remains problematic, and expressibility continues to depend heavily on the basis. On the other hand, Henson et al. [10] investigated also the ring $C(X)$ of real-valued functions on a topological space $X$. Similarly, Eagle et al. [5] associate a Hausdorff compact space $X$ with its $C^{*}$-algebra $C(X)$, and apply continuous logic. This is a promising research direction, however it seems restricted to spaces without any additional structure: what would be the $C^{*}$-algebra associated with a compact graph?
In this study, we take a different approach. The starting point is the notion of a co-valaution on a set $M$, which plays the role of a valuation in first-order logic – that is, an interpretation of free variables. Specifically, a co-valuation is a relation $v\subseteq X\times M$, where $X$ is a finite set (called a context) such that, for $x]_{v}=\{a\in M:(x,a)\in v\}$, the family $v]=\{[x]_{v}:x\in X\}$ is a minimal cover of $M$. A fundamental relation between co-valuations is a consolidation: a covaluation $v\subseteq X\times Y$ is a consolidation of $w\subseteq Y\times M$ – and $w$ is a fragmentation of $v$ – if $[v]$ is refined by $[w]$, and every set $x]_{v}$ is a union of elements of the form $y]_{w}$. The associated consolidation pattern is the relation $\geq\subseteq X\times Y$, defined by $x\geq y$ iff $x]_{v}\supseteq[y]_{w}$, which describes how the consolidation is carried out. A structure $(M,V)$ in the logic of co-valuations is an ordinary first-order structure $M$ together with a family $V$ of co-valuations on the universe of $M$ in contexts drawn from a fixed set of variables, satisfying certain additional requirements. Co-valuations in $V$ are called admissible.
@@
This logic behaves well with respect to ultraproducts: it satisfies Łoś’s theorem (Theorem 5.4), and therefore enjoys compactness (Corollary 5.3). Moreover, with an aid of a duality recently discovered by Bartoš, Bice, and Vignati [2] between certain countable posets, called $\omega$-posets, and second-countable compact $T_{1}$ spaces, for every structure, there exists an elementarily equivalent one whose admissible co-valuations yield a second-countable compact $T_{1}$ topology. We call such structures compact $\omega$-structures. Thus, a topological version of the downward Löwneheim-Skolem theorem holds (Theorem 5.4 ), even though no topology is present in the original setup. Furthemore, the omitting types theorem can be proved (Theorem 5.7), using Baire-category methods in the Polish space $\rm{Mod}(L)$ of all compact $\omega$-structures in signature $L$ endowed with the logic topology. This leads to a natural notion of atomicity for compact $\omega$-structures (Theorem 5.13).
@@
One can show that every compact $T_{1}$ space equipped with closed relations can be naturally presented as a structure in our logic of co-valuations (Corollary 4.15). Since admissible co-valuations form only a basis for the underlying topology, not all topological properties are expressible in the logic. Nevertheless, because in fact we work with covers coming from this basis, rather than the basis itself, notions such as connectedeness, perfectness, or covering dimension are expressible; some others, such as metrizability or chainability follow from omitting a suitable family of types (Section 4.4).
@@
The fact that admissible co-valuations give only a basis is a limitation. However, for many classical topological objects there is a canonical choice of admissible co-valuations. Specifically, this occurs for second-countable compact $T_{1}$ spaces (possibly with additional relations) that arise as Fraïssé limits in a variant of Fraïssé theory suitable for this setting (Section 6). When admissible co-valuations are obtained via a sufficiently nice ($L$-faithful) Fraïssé sequence in an appropriate category, the resulting structure is an atomic model of its theory (Theorem 6.6), and any two such structures are isomorphic (Theorem 6.5). This applies, for example, to the arc, the ordered arc, or the pseudo-arc. Thus, for compact objects arising as Fraïssé limits, there is a canonical way to select the “correct” structure that presents it in the logic of co-valuations.
@@
The paper is organized as follows. Section 3 provides a brief overview of the Bartoš-Bice-Vignati duality between $\omega$-posets and second- countable compact $T_{1}$ spaces as well as some other background facts required later. In Section 4, we introduce the logic of co-valuations, define compact structures, and examine several basic topological properties that are expressible in this logic. Section 5 develops the model theory: ultraproducts, Łoś’s theorem, the downward Löwenheim-Skolem theorem, omitting types, and atomic models. Section 6 presents a variant of Fraïssé theory adapted to the logic of co-valuations, and discusses atomicity of Fraïssé limits along with several examples.
@@ 2. Relations
$(xR^{\prime}y\mbox{ and }x^{\prime}Ry)\rightarrow x=x^{\prime}.$
@@
We say that $A$ is $R$-refined by $B\subseteq Y$ (or that $A$ is refined by $B$, provided that $R$ is clear from the context) if $A^{R}\supseteq B$. Typically, $R\subseteq\mathbb{P}\times\mathbb{P}$ is an ordering relation $\geq$ of a poset $\mathbb{P}$, e.g., the inclusion relation $\supseteq$. In the case that the ordering relation of a poset $\mathbb{P}$ is not explicitely specified, we will refer to it by $\leq$ (or by $\geq$). For $A,B\subseteq\mathbb{P}$, we say that $A$ is an $\leq$-up-set (or up-set) if $A^{\leq}=A$,
@@ 3. Constructing compact spaces from posets
In this section, we present a brief overview of the Bartoš–Bice–Vignati duality connecting certain countable posets and second-countable compact $T_{1}$ spaces in the spirit of Stone duality. It has been developed in [2], to which the reader is referred for a detailed exposition. We provide some proofs to illustrate the role played by the main concepts of this theory.
We begin with some motivation. Let $X$ be a second-countable compact $T_{1}$ space, and let $\mathbb{P}_{n}$, $n\in\mathbb{N}$, be a sequence of finite open covers of $X$ such that every $\mathbb{P}_{n}$ is refined by $\mathbb{P}_{n+1}$, and every open cover of $X$ is refined by some $\mathbb{P}_{n}$. In particular, $\bigcup\mathbb{P}_{n}$ is a basis of $X$ (see Proposition 3.7). For $T_{2}$, i.e., metrizable, compact spaces, the construction of such a sequence, using the Lebesgue number of a finite cover, is straightfoward; the general case is also elementary (see [2, Proposition 1.11]). Let the disjoint union $\mathbb{P}=\bigsqcup\mathbb{P}_{n}$ be ordered by inclusion, i.e., $p\geq q$ iff $p\supseteq q$, where $p\in\mathbb{P}_{m}$, $q\in\mathbb{P}_{n}$, $m\leq n$.
@@
As a matter of fact, we will work with $\omega$-posets of a particularly simple form, where the ordering relation is determined by relations between consequtive levels. Let $\mathbb{P}_{n}$, $n\in\mathbb{N}$, be finite sets, and let $\geq_{n}\subseteq\mathbb{P}_{n}\times\mathbb{P}_{n+1}$, $n\in\mathbb{N}$, be co-surjective relations. Let $\mathbb{P}=\bigsqcup\mathbb{P}_{n}$, and let $\geq$ be the reflexive and transitive closure of $\bigsqcup\geq_{n}$. It is straightforward to verify that the resulting poset $(\mathbb{P},\geq)$ is an $\omega$-poset, referred to as an $\omega$-chain $(\mathbb{P}_{n},\geq_{n})$. For an $\omega$-chain $(\mathbb{P},\geq)$, $\mathbb{P}_{n}$ always denote the levels of $\mathbb{P}$, $\geq^{m}_{n}=\geq\upharpoonright\mathbb{P}_{m}\times\mathbb{P}_{n}$, where $m\leq n$, and $\geq_{n}=\geq^{n}_{n+1}$. We call $(\mathbb{P}_{n},\geq_{n})$ level-injective if all $\geq_{n}$ are co-injective.
@@
Let $A$, $B$ be minimal covers of a set $X$ such that $A$ is refined by $B$. Then $\supseteq\upharpoonright A\times B$ is co-injective.
@@
Fix $p\in A$, and, using minimality of $A$, $x\in p\setminus\bigcup\{p^{\prime}\in A:p^{\prime}\neq p\}$. As $B$ covers $X$, there is $q\in B$ with $x\in q$. As it refines $A$, there is $p^{\prime}\in A$ such that $p^{\prime}\supseteq q$. Clearly, $p^{\prime}\supseteq q$ implies $p^{\prime}=p$ for every $p^{\prime}\in A$. ∎
Proposition 3.5.
Let $\mathbb{P}_{n}$, $n\in\mathbb{N}$, be minimal covers of a set $X$ such that each $\mathbb{P}_{n}$ consolidates $\mathbb{P}_{n+1}$. Then $(\mathbb{P},\geq)=(\mathbb{P}_{n},\supseteq\upharpoonright\mathbb{P}_{n}\times\mathbb{P}_{n+1})$ is a level-injective $\omega$-chain, and, for any $p\in\mathbb{P}_{m}$, and $q\in\mathbb{P}_{n}$, where $m\leq n$, we have
$p\geq q\iff p\supseteq q.$
Thus, slightly abusing notation, we can denote $(\mathbb{P},\geq)$ by $(\bigsqcup\mathbb{P}_{n},\supseteq)$.
Proof.
Level-injectivitiy follows from Proposition 3.4. Let $\supseteq_{n}=\supseteq\upharpoonright\mathbb{P}_{n}\times\mathbb{P}_{n+1}$, and, for $m<n$, denote by $\supseteq^{m}_{n}$ appropriate compositions of $\supseteq_{n}$. We claim that if $\supseteq_{n}^{0}$ is the transitive closure of $\bigcup_{m\leq n}\supseteq_{m}$, then $\supseteq_{n+1}^{0}$ is the transitive closure of $\bigcup_{m\leq n+1}\supseteq_{m}$. Fix $m\in\mathbb{N}$ with $m<n$, and $p\in\mathbb{P}_{m}$, $q\in\mathbb{P}_{n+1}$ with $p\supseteq q$. By minimality of the cover $\mathbb{P}_{n+1}$, there is $x\in q\setminus\bigcup\{q^{\prime}\in\mathbb{P}_{n+1}:q^{\prime}\neq q\}$. Because $\mathbb{P}_{m}$ consolidates $\mathbb{P}_{n}$, there is $q^{\prime}\in\mathbb{P}_{n}$ such that $x\in q^{\prime}\subseteq p$. But then we must have $q\subseteq q^{\prime}$, which immediately implies the claim. ∎
@@
For each $x\in X$, let $d_{x}=\bigcap\{a\in A\cup B:x\in a\}$. As $A$ and $B$ are finite, so is $D=\{d_{x}:x\in X\}$. For each $d\in D$, choose some $x_{d}\in X$ such that $d=d_{x_{d}}$ and denote the set of all the other chosen points by
@@
Also note that if $y\in a\in A$ then $y\neq x_{d}$ for any $d\neq d_{y}$ (because $y=x_{d}$ implies $d_{y}=d_{x_{d}}=d$), so $y\in d_{y}\setminus f_{d_{y}}\subseteq a$. This shows that $a=\bigcup\{c\in C:c\subseteq a\}$, for all $a\in A$, i.e., $C$ fragments $A$. For the same reason, $C$ fragments $B$. ∎
@@
Fix an open $c\subseteq X$, and $x\in c$. Let $C$ be an open cover of $X$ extending $\{c\}$, and such that $x\not\in\bigcup C\setminus\{c\}$. Fix $C^{\prime}\in\mathcal{C}$ that refines $C$, and let $c^{\prime}\in C^{\prime}$ be such that $x\in c^{\prime}$. Then we must have $c^{\prime}\subseteq c$. ∎
@@
For $\omega$-chains $\mathbb{P}$ without minimal elements, e.g., level-injective $\omega$-chains, levels are bands, and caps can be equivalently defined as subsets of $\mathbb{P}$ that are refined by a level.
@@
To prove (1), we use Zorn’s lemma. Let $S_{i}$, $i\in I$, be an infinite chain (under inclusion) of selectors. We show that $S=\bigcap_{i}S_{i}$ is a selector. Let $C$ be a cap. We can assume that $C$ is finite, so there is $c\in C$ such that for $S_{i}\cap C=\{c\}$ for almost all $i\in I$. But $S_{i}$’s form a chain, so, actually, $S_{i}\cap C=\{c\}$ for all $i\in I$.
@@
A basis $\mathbb{P}$ of a topological space $X$ is called a cap-basis if every cover of $X$ consisting of sets from $\mathbb{P}$ is a cap in the poset $\mathbb{P}$ ordered by inclusion.
@@
We prove that $\tau_{\mathbb{P}}$ is a compact topology. Without loss of generality, we can assume that $\mathsf{S}$$\mathbb{P}$, and so $\mathbb{P}$, is infinite. Fix a cap $C$, and a band $B$ that refines $C$. Clearly, $B^{\in}$ is a band in $\mathbb{P}^{\in}$, i.e., $C^{\in}$ is a cap. Fix $S\in{\mathsf{S}}\mathbb{P}$. As $S$ is infinite, $B$ is finite, and $\mathbb{P}$ is an $\omega$-poset, there must exist $s\in S$, and $b\in B$ such that $s\leq b$. Because $B$ refines $C$, there is $c\in C$ such that $b\leq c$. But $S$ is an up-set, so $c\in S$, i.e., $S\in c^{\in}$. As $S$ was arbitrary, we get that $C^{\in}$ covers $\mathsf{S}$$\mathbb{P}$.
Suppose now that $C\subseteq\mathbb{P}$ is not a cap. In other words, $\mathbb{P}\setminus C$ is a selector, so it contains a minimal selector $S$. But then $c\not\in S$, i.e., $S\not\in c^{\in}$ for every $c\in C$, and $C^{\in}$ does not cover $\mathsf{S}$$\mathbb{P}$. Therefore covers of $X$ by sets from $\mathbb{P}^{\in}$ are given by caps in $\mathbb{P}$. In particular, $\mathbb{P}^{\in}$ is a cap-basis, and $\tau_{\mathbb{P}}$ is compact. ∎
@@
If $\mathbb{P}$ is a level-injective $\omega$-chain, then every level $\mathbb{P}_{n}^{\in}$ is a minimal $\tau_{\mathbb{P}}$-open cover of $\mathsf{S}$$\mathbb{P}$, and every $\tau_{\mathbb{P}}$-open cover is refined by some $\mathbb{P}_{n}^{\in}$. In particular, $^{\in}\neq\emptyset, pp\in\mathbb{P}$.
@@
The facts that every $\mathbb{P}_{n}^{\in}$ is a $\tau_{\mathbb{P}}$-open cover of $\mathsf{S}$$\mathbb{P}$, and every $\tau_{\mathbb{P}}$-open cover is refined by some $\mathbb{P}_{n}^{\in}$, directly follow from Remark 3.8 and Proposition 3.12. It remains to show minimality of $\mathbb{P}_{n}^{\in}$’s.
@@
Proof.
First we observe, arguing as in the introduction to Section 3, that every $x^{\in}$ is indeed a minimal selector in $\mathbb{P}$. As $X$ is $T_{1}$, for any distinct $x,x^{\prime}\in X$, there is $p\in\mathbb{P}$ such that $x\in p$, $x^{\prime}\not\in p$, i.e., the mapping $x\mapsto x^{\in}$ is injective. Finally, we note that, for every $p\in\mathbb{P}$, $x\in p$ iff $x^{\in}\in p^{\in}$, which implies that it is continuous, and its image is dense in $\mathsf{S}$$\mathbb{P}$. Thus, it is onto $\mathsf{S}$$\mathbb{P}$, i.e., a homeomorphism between $X$ and $\mathsf{S}$$\mathbb{P}$. ∎
@@
Note that $x\in p$ iff $x^{\in}\in p$, so the homeomorphism $x\mapsto x^{\in}$ maps $p\in\mathbb{P}$ onto $p^{\in}$.
@@
Let $\mathbb{P}$ be a level-injective $\omega$-chain. Then $x\geq y$ iff $x^{\in}\supseteq y^{\in}$, for $x\in\mathbb{P}_{m}$, $y\in\mathbb{P}_{n}$, where $m\leq n$.